Parth (parthkohli):

A 100W bulb and a 25W bulb are designed for the same voltage. They have filaments of the same length and material. The ratio of the diameter of the 100W bulb to the 25W bulb is?

Parth (parthkohli):

Erm, so...\[\dfrac{V^2}{R_1}= 100\]\[\dfrac{V^2}{R_2}=25\]So\[\dfrac{R_1}{R_2}=\dfrac{1}{4}\]

Parth (parthkohli):

\[\dfrac{A_1}{A_2} = \dfrac{1}{4}\]\[\dfrac{r_1}{r_2} = \dfrac{1}{2}\]

Parth (parthkohli):

Whoops, I meant\[\dfrac{r_1}{r_2} = 2\]

Parth (parthkohli):

The EMF of a cell is \(\mathcal E\) and its internal resistance is \(r\). Its terminals are connected to a resistance \(R\). The potential difference between the terminals is \(1.6 V\) for \(R = 4 \Omega\), and \(1.8 V\) for \(R = 9 \Omega\). Find \(\mathcal E\) and \(r\).

Parth (parthkohli):



Parth (parthkohli):

\[i = \dfrac{\mathcal E}{R + r}\]

Parth (parthkohli):

\[- \dfrac{\mathcal E}{4 +r}\cdot 4 - \dfrac{\mathcal E}{4 + r}\cdot r=1.6\]